# Just Trial Once: Ongoing Causal Validation of Machine Learning Models

### v1

@@ 4 Implications for pre-trial design and post-trial analysis
Recommendation: Conduct trials with multiple models. Our results suggest the utility of trialing multiple models in a cluster RCT that vary in their outputs on different patient populations and which exhibit a range of reasonable performance characteristics. First of all, doing so gives more flexibility to estimate the value of new models if there are sizeable populations where trialed models raise different outputs. Second, falsification of our main assumptions (Assumptions 3.1 and 3.2) can be done using data from patient populations where the outputs of different models agree.
@@
Recommendation: Improved accuracy is not sufficient for deployment. Given a set of models or policies included in a trial, it may be tempting to conclude that an updated model or policy that is more accurate (on average) should be deployed. However, this conclusion may be flawed given the goal of improving patient outcomes. For instance, a model that is more accurate, but achieves higher accuracy by sacrificing performance on some important subpopulation, may ultimately lead to worse outcomes. Indeed, in Section 5 we give a simple simulated example where the optimal model is not the model with the best performance.
@@ 5 Simulation Study
We now describe a simple simulated example, inspired by Example 1, that demonstrates the results derived in Section 3 and how our proposed method allows for more robust comparisons between models. We consider machine learning models that alert clinicians to the near-term onset of some disease, denoted $O\in\{0,1\}$. We simulate a cluster RCT with three arms: A control arm, denoted as a policy $\pi_{0}$ that never alerts, and two arms where models $\pi_{1}$ and $\pi_{2}$ are deployed, respectively. For simplicity, we consider model performance $M$ to be the ground-truth accuracy in predicting disease onset. Our outcome of interest, $Y\in\{0,1\}$, is patient survival, and $X$ represents a baseline health characteristics. For the sake of an interpretable simulation, $X$ takes on values uniformly in $\{0,1,2,3\}$, corresponding to decreasing likelihood of disease onset. The trialed models are defined as $\pi_{1}=\mathbf{1}\{X=1\}$ and $\pi_{2}=\mathbf{1}\{X\in\{2,3\}\}$.
Table 1 gives expected values of $O$ and $Y(a,m)$ over all possible values of $X$, satisfying Assumptions 3.1, 3.2 and 3.3. In this simulation, patients with $X=0$ are the most likely to develop disease, yet raising an alert on these patients only increases their survival rate by $0.1$, while raising an alert on other cases increases survival rate by $0.5$. This simulated scenario captures the real-world intuition that clinicians may already be paying attention to these “obvious” cases, such that raising an alert does little to improve outcomes.
Figure 4: Bar graph showing accuracy at predicting disease, and expected patient survival rates, of different simulated alerting models. Bars denote ground-truth values, and blue intervals denote estimated bounds using our approach. Because $\pi_{e0}$ and $\pi_{e1}$ were not trialed in the RCT, their bounds are wider. Despite $\pi_{e1}$ having the greatest raw accuracy, $\pi_{e0}$ has the most positive effect on patient survival rates.
We consider the effect of deploying two new models, $\pi_{e0}=\mathbf{1}\{X\in\{1,2,3\}\}$ and $\pi_{e1}=\mathbf{1}\{X\in\{0,1\}\}$, whose accuracy at predicting onset is computable using Table 1. We then estimate bounds on $\mathbb{E}[Y(A=\pi,M=f_{M}(\pi))]$ for each model (including confidence intervals to incorporate finite-sample uncertainty, as described under Assumption 3.3) using data from a simulation with $n=5,000$. Fig. 4 shows the simulation results and illustrates that the most accurate model is not always the best: $\pi_{e1}$ has the greatest accuracy in predicting the onset of disease, but $\pi_{e0}$ has the largest causal impact on patient outcomes. This reversal occurs since $\pi_{e0}$ raises alerts for patients who stand to benefit the most, whereas $\pi_{e1}$ tends to alert for patients who have little to gain from them. Moreover, our bounds reflect greater confidence in the (positive) impact of $\pi_{e0}$, since the lower bound for patient outcomes under $\pi_{e0}$ is greater than patient outcomes under all trialed models. Python code implementing this simulation study, implementing the estimation procedure proposed in Assumption 3.3, and for generating Fig. 4 is publicly available online at: https://github.com/jacobmchen/just_trial_once.
6 Conclusion and Limitations
In this paper, we discussed methods for estimating the causal impact of new or updated ML and AI models not previously trialed in an RCT. Under the important considerations that ML predictions are deterministic and that clinician trust in ML models play a role in determining their impacts on patient outcomes, we demonstrated how one could estimate lower and upper bounds on a new model. Given the possibility that key assumptions for employing our methods may not be fulfilled, we proposed simple strategies for testing and falsifying them. Finally, we concluded with a simulation study to illustrate the application of our method and to highlight its benefits when selecting among model updates and considering the causal impact of new models.
@@ 6 Conclusion and Limitations
$\mathbb{E}[Y(a,m)]$ $0.3+0.5a+0.1am$ $0.2+0.5a+0.1am$
Table 1: A table summarizing the probability of developing disease ($\mathbb{E}[O]$) and the expected survival rate ($\mathbb{E}[Y(a,m)]$) for the simulation in Section 5.
@@
However, our work is not without limitations: First, our derived bounds are naturally pessimistic, and, while they cannot be tightened without further assumptions (Assumption 3.3), some additional assumptions may be warranted in some cases. For instance, we implicitly assume that “anything can happen” when a model raises an alert on patients who never received alerts in the past. One could instead assume that alerts are not harmful (except for their impact on performance), or that they are not harmful for a specific patient if the alert is correct. Second, we assume that model performance can be summarized in a single real number, but a more complex representation (e.g., involving subgroup-specific performance) may be warranted in some applications. Finally, a core limitation of our approach is that we still require an RCT. Using RCT data comes with many benefits: It allows for greater confidence that core assumptions (e.g., randomization of policies) hold by design, and even allows for checking (in some cases) the core assumptions we make in this paper, as we have shown (Assumptions 3.1 and 3.2). However, observational studies (e.g., pre-post studies of model deployments) are often easier to run in practice. Our hope is that this work can serve as a springboard towards increasing the utility, and thus adoption, of RCTs for ML models deployed in high-risk settings.
Appendix
@@ Appendix A Brief Causal Inference Overview
First, independencies implied in the full data distribution $p({\bf V})$ can be read off from $\mathcal{G}$ using the d-separation criterion [Pearl, 2014]. In our proofs, we assume faithfulness [Spirtes et al., 2001], which states that any independencies implied by the data distribution must be reflected in $\mathcal{G}$ via d-separation and vice versa. Formally, $A\perp\!\!\!\perp B\mid{\bf C}$ in $p({\bf V})$ if and only if $A\perp\!\!\!\perp_{\text{d-sep}}B\mid{\bf C}$, where $A$ and $B$ are variables and ${\bf C}$ is a set of variables that can be empty.
Next, we follow [Richardson and Robins, 2013] to convert a DAG $\mathcal{G}$, given an intervened on variable $W$, to a single-world intervention graph (SWIG) using a node-splitting transformation as follows: (1) replace all children of $W$, $V$, with the potential outcome random variable $V(w)$, (2) add the intervened on variable $w$ into $\mathcal{G}$ as a new vertex, and (3) change all outgoing edges from $W$ to originate from $w$ instead. This new SWIG is denoted by $\mathcal{G}(w)$. It is possible to intervene on multiple variables in $\mathcal{G}$ by repeatedly applying the node-splitting transformation described above. Fig. 5 shows the SWIG $\mathcal{G}(a,m)$, the SWIG containing the potential outcome random variable $Y(a,m)$ corresponding to the target estimand $\mathbb{E}[Y(A=\pi_{e},M=f_{M}(\pi_{e}))]$.
In order to connect potential outcome distributions to observed distributions, [Malinsky et al., 2019] propose a set of three rules known as the potential outcome calculus (po-calculus). Most salient to our proofs is Rule 2 of po-calculus, which states that $p(V(w)\mid{\bf X})=p(V\mid{\bf X},W=w)$ if $V(w)\perp\!\!\!\perp W\mid{\bf X}$ in $\mathcal{G}(w)$, where ${\bf X}$ is any set of random variables in $\mathcal{G}(w)$, including the empty set. Rule 2 is also referred to as the conditional ignorability assumption commonly used in causal inference, which states that a potential outcome of interest is independent of a treatment of interest conditional on a sufficiently rich set of covariates ${\bf X}$. The benefit of using Rule 2 of po-calculus is that it allows us to use d-separation in a SWIG to directly verify whether conditional ignorability holds. For instance, note that $Y(a,m)\perp\!\!\!\perp A,M\mid X$ in the SWIG shown in Fig. 5. Rule 2 of po-calculus thus allows us to conclude that $p(Y(a,m)\mid X)=p(Y\mid A=a,M=m,X)$.
@@
Appendix B Proofs

### v2

@@ 4 Recommendations For Pre-trial Design
Recommendation: Conduct trials with multiple models. Our results suggest the utility of trialing multiple models in a cluster RCT that vary in their outputs on different patient populations and which exhibit a range of reasonable performance characteristics. First of all, doing so gives more flexibility to estimate the effect of new models if there are sizeable populations where trialed models raise different outputs. Second, falsification of our main assumptions (Assumptions 3.1 and 3.2) can be done using data from patient populations where the outputs of different models agree.
@@
Recommendation: Use previous trial data to inform deployment of new models. Given a set of models or policies included in a trial, it may be tempting to conclude that an updated model that is more accurate (on average) should be deployed. However, this conclusion may be flawed given the goal of improving patient outcomes. For instance, a model that is more accurate, but achieves higher accuracy by sacrificing performance on some important subpopulation, may ultimately lead to worse outcomes. Indeed, in Section 5, we give a simple simulated example where the optimal model is not the model with the best performance. Given the potential impacts of deploying ML models in practice, it is paramount that we use previous trial data to draw inference on outcomes of interest prior to deploying a new model.
@@ 5 Simulation Study
We now describe a simple simulated example, inspired by Example 1, that demonstrates the results derived in Section 3 and how our proposed method allows for more robust comparisons between models. We consider machine learning models that alert clinicians to the near-term onset of some disease, denoted by $O\in\{0,1\}$. We simulate a cluster RCT with three arms: A control arm, denoted as a policy $\pi_{0}$ that never alerts, and two arms where models $\pi_{1}$ and $\pi_{2}$ are deployed, respectively. For simplicity, we consider model performance $M$ to be the ground-truth accuracy in predicting disease onset. Our outcome of interest, denoted by $Y\in\{0,1\}$, is patient survival, and $X$ represents a baseline health characteristic. For the sake of an interpretable simulation, $X$ takes on values uniformly in $\{0,1,2,3\}$, and $O$ and $Y$ are Bernoulli random variables. The probability of $O=1$ depends only on $X$ whereas the probability of $Y=1$ depends on $X$, $A$, and $M$. The trialed models are defined as $\pi_{1}=\mathbf{1}\{X=1\}$ and $\pi_{2}=\mathbf{1}\{X\in\{2,3\}\}$.
Figure 4: Bar graph giving a visual representation of the data generating process in the simulation study. The bars on the left depict the likelihood of disease onset for varying levels of baseline health $X$, and the bars on the right depict the likelihood of survival, assuming model performance is fixed at $m=0.5$. The arrows above the bars on the right show how survival rates change for each level of $X$ when patients receive an alert from an alerting model.
Table 1: Probability of developing disease ($\mathbb{E}[O]$) and expected survival rate ($\mathbb{E}[Y(a,m)]$) for the simulation study.
@@
$\mathbb{E}[Y(a,m)]$ $0.46+((1+m)/2)\cdot 0.48a$ $0.55+((1+m)/2)\cdot 0.4a$
Table 1 gives expected values of $O$ and $Y(a,m)$ over all possible values of $X$, satisfying Assumptions 2.1, 3.2 and 3.3; Fig. 4 visualizes the same information for fixed model performance. The aggregate patterns reflect the following process: First, we assume that all patients who do not develop disease will survive. Then, among the sickest individuals ($X=0$), the survival rate among patients with disease is 40% without alerts [^10] To align this explanation with Table 1, note that all of the 10% without disease and $40\%$ of the $90\%$ with disease survive, yielding the overall survival rate of 46% without alerts. and between 50-60% with alerts, as influenced by $m$, where 50% corresponds to a model with zero accuracy, and 60% is achieved by a model with perfect accuracy. These patterns reflect the intuition that survival is improved (when alerts are raised) under a model perceived to be more accurate. However, the survival rate (among those with disease) is much improved for other groups, going from 10% without alerts to 50-90% with alerts [^11] For explicit computations of survival rates with and without alerts, see Appendix C.. Since the actual prevalence of disease varies across these groups, and since alerts only help those with disease, the overall causal effect is largest among those where $X=1$. Appendix C describes and explains this data generating process in detail.
Figure 5: Bar graph showing accuracy at predicting disease and expected patient survival rates of different simulated alerting models. Bars denote ground-truth values, and blue intervals denote estimated bounds using our approach. Because $\pi_{e0}$ and $\pi_{e1}$ were not trialed in the RCT, their bounds are wider. Despite $\pi_{e1}$ having the greatest raw accuracy, $\pi_{e0}$ has the most positive effect on patient survival rates.
We consider the effect of deploying two new models, $\pi_{e0}=\mathbf{1}\{X\in\{1,2,3\}\}$ and $\pi_{e1}=\mathbf{1}\{X\in\{0,1\}\}$, whose accuracies at predicting onset are computable using Table 1. We then estimate bounds on $\mathbb{E}[Y(A=\pi,M=f_{M}(\pi))]$ for each model (including confidence intervals to incorporate finite-sample uncertainty, as described under Remark 3.1) using data from a simulation with $n=5,000$. Fig. 5 shows the simulation results and illustrates that the most accurate model is not always the best: $\pi_{e1}$ has the greatest accuracy in predicting the onset of disease, but $\pi_{e0}$ has the largest causal impact on patient outcomes. This reversal occurs since $\pi_{e0}$ raises alerts for patients who stand to benefit the most, whereas $\pi_{e1}$ tends to alert for patients who have little to gain from them. Moreover, our bounds reflect greater confidence in the (positive) impact of $\pi_{e0}$, since the lower bound for patient outcomes under $\pi_{e0}$ is greater than patient outcomes under all trialed models. Python code implementing this simulation study, implementing the estimation procedure proposed in Remark 3.1, and for generating Figs. 4 and 5 are publicly available online at: https://github.com/jacobmchen/just_trial_once.
6 Conclusion and Limitations
In this paper, we discussed methods for estimating the causal impact of new or updated ML and AI models not previously trialed in an RCT. Under the important considerations that ML predictions are deterministic and that clinician trust in ML models play a role in determining their impacts on patient outcomes, we demonstrated how one could estimate lower and upper bounds on the effect of deploying a new model. We further proved tightness of our proposed bounds (Remark 3.1) and gave inverse probability weighted-style estimators for them (Remark 3.1). Given the possibility that key assumptions for employing our methods may not be fulfilled, we also proposed simple strategies for testing and falsifying them. Finally, we concluded with a simulation study to illustrate the application of our method and to highlight its benefits when selecting among model updates.
@@ 6 Conclusion and Limitations
However, our work is not without limitations: First, our derived bounds are naturally pessimistic, and, while they cannot be tightened without further assumptions (Remark 3.1), some additional assumptions may be warranted in certain cases. For instance, we implicitly assume that “anything can happen” when a model raises an alert on patients who never received alerts in the past. One could instead assume that alerts are not harmful (except for their impact on performance), or that they are not harmful for a specific patient if the alert is correct. Second, we assume that model performance can be summarized in a single real number, but a more complex representation (e.g., involving subgroup-specific performance, performance adaptation over time, and clinician experience) may be warranted in some applications. Finally, a core limitation of our approach is that we still require an RCT. Using RCT data comes with many benefits: It allows for greater confidence that core assumptions (e.g., randomization of policies) hold by design, and even allows for checking (in some cases) the core assumptions we make in this paper, as we have shown (Assumptions 3.1 and 3.2). However, observational studies (e.g., pre-post studies of model deployments) are often easier to run in practice. Our hope is that this work can serve as a springboard towards increasing the utility, and thus adoption, of RCTs for ML models deployed in high-risk settings.
Acknowledgements.
The authors are grateful for the helpful feedback of anonymous reviewers as well as fruitful discussions with Jonathan Zhang, Erik Skalnes, and Trung Phung.
Supplementary Material
@@ Appendix A Brief Causal Inference Overview
First, independencies implied in the full data distribution $P({\bf V})$ can be read off from $\mathcal{G}$ using the d-separation criterion [Pearl, 2014]. Next, we follow [Richardson and Robins, 2013] to convert a DAG $\mathcal{G}$, given an intervened on variable $W$, to a single-world intervention graph (SWIG) using a node-splitting transformation as follows: (1) replace all children of $W$ (here, $V$) with the potential outcome random variable $V(w)$, (2) add the intervened on variable $w$ into $\mathcal{G}$ as a new vertex, and (3) change all outgoing edges from $W$ to originate from $w$ instead. This new SWIG is denoted by $\mathcal{G}(w)$. It is possible to intervene on multiple variables in $\mathcal{G}$ by repeatedly applying the node-splitting transformation described above. Fig. 6 shows the SWIG $\mathcal{G}(a,m)$, the SWIG containing the potential outcome random variable $Y(a,m)$ corresponding to the target estimand $\mathbb{E}[Y(A=\pi_{e},M=f_{M}(\pi_{e}))]$.
In order to connect potential outcome distributions to observed distributions, [Malinsky et al., 2019] propose a set of three rules known as the potential outcome calculus (po-calculus). Most salient to our proofs is Rule 2 of po-calculus, which states that $P(V(w)\mid{\bf X})=P(V\mid{\bf X},W=w)$ if $V(w)\perp\!\!\!\perp W\mid{\bf X}$ in $\mathcal{G}(w)$, where ${\bf X}$ is any set of random variables in $\mathcal{G}(w)$, including the empty set. Rule 2 is also referred to as the conditional ignorability assumption commonly used in causal inference, which states that a potential outcome of interest is independent of a treatment of interest conditional on a sufficiently rich set of covariates ${\bf X}$. The benefit of using Rule 2 of po-calculus is that it allows us to use d-separation in a SWIG to directly verify whether conditional ignorability holds given a graph. For instance, note that $Y(a,m)\perp\!\!\!\perp A,M\mid X$ in the SWIG shown in Fig. 6. Rule 2 of po-calculus thus allows us to conclude that $P(Y(a,m)\mid X)=P(Y\mid A=a,M=m,X)$.
@@
Appendix B Algorithmic Viewpoint of Computing Bounds
Here, we give an algorithmic construction of the lower and upper bounds presented in Remark 3.1. First, we define sub-algorithms in Algorithm 1 that compute the lower and upper bounds conditional on a value of $x\in\mathcal{X}$. Next, we define the algorithm that constructs the complete bounds in Algorithm 2 by iterating through all possible values of $x\in\mathcal{X}$ and computing the weighted average of the bounds for each $x$.
Algorithm 1 Sub-Algorithms for computing lower and upper bounds conditional on $X$.
1:Input: $x$
2:Output: Tight lower bound on $Y$ conditional on $X=x$
3:function Conditional_Lower($x$)
4:     if $\pi_{e}(x)\neq a_{0}$ then
5:         if $\tilde{\mathbf{\Pi}}^{e}_{\leq}(x)\neq\varnothing$ then return $\mathbb{E}[Y\mid X=x,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(x)]$
6:         else if $\tilde{\mathbf{\Pi}}^{e}_{\leq}(x)=\varnothing$ then return $Y_{min}$
7:         end if
8:     else if $\pi_{e}(x)=a_{0}$ then
9:         if $\mathbf{\Pi}^{e}(x)\neq\varnothing$ then return $\mathbb{E}[Y\mid X=x,\Pi\in\mathbf{\Pi}^{e}(x)]$
10:         else if $\mathbf{\Pi}^{e}(x)=\varnothing$ then return $Y_{min}$
11:         end if
12:     end if
13:end function
14:
15:Input: $x$
16:Output: Tight upper bound on $Y$ conditional on $X=x$
17:function Conditional_Upper($x$)
18:     if $\pi_{e}(x)\neq a_{0}$ then
19:         if $\tilde{\mathbf{\Pi}}^{e}_{\geq}(x)\neq\varnothing$ then return $\mathbb{E}[Y\mid X=x,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\geq}(x)]$
20:         else if $\tilde{\mathbf{\Pi}}^{e}_{\geq}(x)=\varnothing$ then return $Y_{max}$
21:         end if
22:     else if $\pi_{e}(x)=a_{0}$ then
23:         if $\mathbf{\Pi}^{e}(x)\neq\varnothing$ then return $\mathbb{E}[Y\mid X=x,\Pi\in\mathbf{\Pi}^{e}(x)]$
24:         else if $\mathbf{\Pi}^{e}(x)=\varnothing$ then return $Y_{max}$
25:         end if
26:     end if
27:end function
Algorithm 2 Algorithm for computing lower and upper bounds for $\mathbb{E}[Y(A=\pi_{e},M=f_{M}(\pi_{e})]$.
1:$lower\_bound\leftarrow 0$
2:$upper\_bound\leftarrow 0$
3:for $x\in\mathcal{X}$ do
4:     $lower\_bound\leftarrow lower\_bound+\textsc{Conditional\_Lower}(x)\times P(X=x)$
5:     $upper\_bound\leftarrow upper\_bound+\textsc{Conditional\_Upper}(x)\times P(X=x)$
6:end for
7:return $(lower\_bound,upper\_bound)$
Appendix C Data Generating Process of Simulation Study
Here, we define in detail the data generating process for the simulation study in Section 5. The trialed models are defined as $\pi_{0}=0$, $\pi_{1}=\mathbf{1}\{X=1\}$, and $\pi_{2}=\mathbf{1}\{X\in\{2,3\}\}$. We use $\text{acc}(\pi_{i})$ to denote the accuracy of policy $\pi_{i}$ at predicting disease $O$. One can verify that this data generating process obeys the structural causal model given in Assumption 2.1.
$X$ $\sim\text{discrete uniform}[0,3]$
$D$ $\sim\text{discrete uniform}[0,2]$
$\Pi$ $=\pi_{D}$
$A$ $=\pi_{D}(X)$
$O$ $\sim\text{Bernoulli}(\mathbf{1}\{X=0\}0.9+\mathbf{1}\{X=1\}0.7+\mathbf{1}\{X=2\}0.6+\mathbf{1}\{X=3\}0.5)$
$M$ $=\text{acc}(\Pi)$
$Y$ $\sim\text{Bernoulli}(\mathbf{1}\{X=0\}(0.46+((1+m)/2)\cdot 0.18A)+\mathbf{1}\{X=1\}(0.37+((1+m)/2)\cdot 0.56A)$
$\quad\quad+\mathbf{1}\{X=2\}(0.46+((1+m)/2)\cdot 0.48A)+\mathbf{1}\{X=3\}(0.55+((1+m)/2)\cdot 0.4A)$
We now explain in detail how the probabilities of the survival rate $Y$ are computed. First, conditional on patients who have the disease ($O=1$), the survival rates without and with alerts are as follows:
Table 2: Probability of survival ($\mathbb{E}[Y(a,m)]$) under no alerts ($\mathbb{E}[Y(0,m)]$) and alerts ($\mathbb{E}[Y(1,m)]$) for the simulation study.
$X=0$ $X=1$
$\mathbb{E}[Y(0,m)\mid O=1]$ $0.4$ $0.1$
$\mathbb{E}[Y(1,m)\mid O=1]$ $0.4+((1+m)/2)\cdot 0.2$ $0.1+((1+m)/2)\cdot 0.8$
$X=2$ $X=3$
$\mathbb{E}[Y(0,m)\mid O=1]$ $0.1$ $0.1$
$\mathbb{E}[Y(1,m)\mid O=1]$ $0.1+((1+m)/2)\cdot 0.8$ $0.1+((1+m)/2)\cdot 0.8$
When patients do not have the disease ($O=0$), their probability of survival is $1$, that is $\mathbb{E}[Y(a,m)\mid O=0]=1$ for all values of $A$ and $M$. The assumption that lower model performance has a weakening effect on the positive effect of alerting is implicit in Table 2. For patients with $X=0$ and $O=1$, the survival probability with alerts range from $0.5$ when accuracy is 00 to $0.6$ when accuracy is $1$. For all other patients, the survival probability with alerts range from $0.5$ when accuracy is 00 to $0.9$ when accuracy is $1$.
Now, to compute the probabilities of survival given in Table 1, we apply the law of total probability – $\mathbb{E}[Y(a,m)]=\mathbb{E}[Y(a,m)\mid O=0]p(O=0)+\mathbb{E}[Y(a,m)\mid O=1]p(O=1)$ – for each value of $X$. We give this explicit computation in the following table.
Table 3: Probability of survival ($\mathbb{E}[Y(a,m)]$) computed explicitly when summed over the likelihood of developing the disease ($O=1$).
$X=0$ $X=1$
$\mathbb{E}[Y(a,m)]$ $1\cdot 0.1+(0.4+((1+m)/2)0.2a)\cdot 0.9$ $1\cdot 0.3+(0.1+((1+m)/2)0.8a)\cdot 0.7$
$X=2$ $X=3$
$\mathbb{E}[Y(a,m)]$ $1\cdot 0.4+(0.1+((1+m)/2)0.8a)\cdot 0.6$ $1\cdot 0.5+(0.1+((1+m)/2)0.8a)\cdot 0.5$
After simplifying the expressions above, we get the probabilities of survival shown in Table 1. Note that the dependence of $Y$ on $O$ does not violate Assumption 2.1 and does not change Fig. 2 because $O$ is an unobserved variable on the directed path from $X$ to $Y$.
Appendix D Proofs